Rosa has designed a pulley system for a school science project. She has determined that as the coefficient of friction increases, the efficiency of the pulley decreases. With a coefficient of friction of 0.2, the efficiency of the pulley is 75%. What would you expect the efficiency to be if the coefficient of friction increased to 0.3?
Answer by Theo(13342)

the formula for inverse variation is y = k/x
when x = .2, y = 75%
plug those values into the formula to find k.
y = k/x becomes 75 = k/.2
solve for k to get k = .2 * 75 = 15
you have k = 15.
now that you found k, you can use it to solve for y when x = .3
the new equation of y = k/x becomes y = 15 / .3
this turns out to be y = .50
when x = .2, y = 75
when x = .3, y = 50
the 2 pulleys are inversely proportional by the constant factor of k.
x goes from .2 to .3
y goes from 75 to 50